A chemist measured the amount of magnesium oxide produced during an experiment . She finds that 8.7 g of magnesium oxide is produced .calculste the number of moles of magnesium oxide produced

Be sure your answer has the correct number of significant digits
Answer #1

The molar mass of Magnesium oxide (MgO) is 24.3+16.0=40.3 g/mol

The number of moles of MgO is the ratio of the mass of MgO to its molar mass

\text {Number of moles of MgO} = \frac {\text {Mass of MgO}}{\text {Molar mass of MgO}} = \frac {8.7 \: g}{40.3 \: g/mol} = 0.22 \: mol